In this exercise they want you to find the percent, parts out of 100, decimal and/or fraction.
#6 is parts out of 100=10.5, find percent, decimal and fraction
#8 is decimal=0.80, find percent, parts out of 100 and fraction
#10 is fraction=3/20, find percent, parts out of 100 and decimal

Answer by checkley77(12844)

#6 is parts out of 100=10.5, find percent, decimal and fraction
10.5/100=.105 DECIMAL
.105=10.5% PERCENT
10.5/100=21/200
--------------------------------------------
#8 is decimal=0.80, find percent, parts out of 100 and fraction
.8=80% PERCENT
.8/100
.8/100=4/500
#10 is fraction=3/20, find percent, parts out of 100 and decimal
3/20=.15 DECIMAL
.15=15% PERCENT
3/20=3*5/20*5=15/100